OpenStudy (anonymous):

7000 dollars is invested in a bank account at an interest rate of 9 per cent per year, compounded continuously. Meanwhile, 20000 dollars is invested in a bank account at an interest rate of 3 percent compounded annually. To the nearest year, when will the two accounts have the same balance?

OpenStudy (anonymous):

x(1+i)^n is the compund interest formula so 7000(1+0.09)^n=20000(1+0.03)^n 7(1.09)^n=20(1.03)^n 1.09^n-(20(1.03)^n)/7 I don't know where to go from here.

OpenStudy (turingtest):

continuously compounded the formula is \[Pe^{tr}=7000e^{0.9t}\]we want to know when that equals our other formula:\[P(1+{r \over n})^{nt}=2000(1+0.03)^{t}\]because n=1 once per year Set equal and solve for t.